OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)

OpenStudy (anonymous):

first find the slope

OpenStudy (anonymous):

then plug the numbers in to slope intercept form of \[y-y_1 = m(x-x_1) \] then turn that into\[ y =mx+b \] form remember m is the slope

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 7}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ 5}}\quad ,&{\color{blue}{ 4}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)